# What Makes LLM Agent Simulations Useful for Policy? Insights From an Iterative Design Engagement in Emergency Preparedness

### v1

@@ 4.1. Iterative Design and Participants
We conducted an iterative design engagement with a university’s emergency preparedness and response team between May 2024 and August 2025 — a span of 16 months. This collaboration was structured around flexible meetings convened as opportunities and needs emerged. While the broader team included many stakeholders, our design work primarily involved five core members who were directly responsible for preparedness and crisis communication: a Senior Director for Disaster Recovery and Business Continuity Services (P1), a Senior Director for Reputation and Issues Management (P2), a Director for Social Media (P3), and two Emergency Preparedness and Response Specialists (P4, P5). Together, these individuals represented both policy-level decision makers and operational specialists, allowing our process to bridge institutional strategy with on-the-ground practice.
@@ 4.2. Data Collection and Analysis
We analyzed our qualitative data as follows. The first author transcribed and coded the meeting transcripts, the interviews, the policy implementation proposal session, and relevant email exchanges. Two other researchers independently reviewed the coded transcripts and provided feedback. The first author began by open coding one transcript, then discussed the emerging codes with the other coders to align on coding granularity and analytic focus. We then applied open coding across the full set of transcripts and emails, capturing both immediate design concerns and broader reflections about policy relevance (Strauss, 1987). Together, we refined the codes into higher-level categories and conducted a thematic analysis to surface recurrent patterns (Braun and Clarke, 2006). We resolved disagreements through discussion until consensus was reached. These qualitative analyses provided the foundation for the detailed accounts of our design iterations ans findings presented in the following sections.
@@ 5. Design Iterations
We developed our simulation system through multiple iterations throughout the 16-month engagement, with each version incorporating feedback from policymakers and addressing specific design challenges (Fig. 2). In this section, we describe this iterative process that revealed how LLM agent simulations evolve from academic demonstrations with abstract potential to practically useful tools for policy implementation.
Figure 2. Iterative design process. Across 16 months (May 2024–Aug 2025), we progressed from preparation interviews through five simulation iterations and an in-situ observation. Each phase introduced greater scale and realism (from 100 to 13,000 agents, adding roles, bottlenecks, and family dynamics) and was shaped by policymakers’ feedback. Early iterations surfaced distrust and missing realism, later iterations built credibility and training value, and by the final iteration, simulations informed adopted protocols, feasibility assessments, and official after-action reports.:
@@ 5.1. Preparation interviews: Capturing Needs and Institutional Realities
To ground and initiate our design process, we conducted three formative interviews during the first five months, complemented by co-creation activities to develop stakeholder and process maps (Newcombe, 2003; Hunt, 1996).
The first interview, with P1, P2, and P3, followed a semi-structured format and focused on current emergency preparedness practices. We asked them to describe prior experiences, identify challenges in existing workflows, and discuss possible barriers to adopting new technologies. We also invited reflections on possible roles for simulation in their work.
The second interview, conducted with P1, examined the policy-making workflow in detail. We asked P1 to walk us through the sequence of activities involved in developing and implementing emergency policies, grounded in concrete prior experiences. During the session, we co-created a stakeholder map and a process map, documenting actors, responsibilities, and information flows (Fig. A1). After the session, we refined the drafts and invited P1 to verify their accuracy.
The third interview, conducted with P3, investigated daily communication practices. We asked P3 to describe concrete routines for monitoring and responding to online interactions during emergencies, and probed specific cases where social media shaped people’s perception or altered information exchange.
All three interviews followed a semi-structured format, combining prepared questions with targeted follow-ups to capture both routine activities and situational variations. The finalized stakeholder and process maps are included in the Appendix (Figs.  A2, A3, A4, and A5).
@@ 5.2. First Iteration: Simulating Communication and Misinformation Dynamics
5.2.1. Simulation Goals and Design
Based on initial interviews with policymakers, we identified communication challenges as a primary concern during emergencies. P1, P2 and P3 expressed interest in understanding how official announcements might be misinterpreted and how misinformation could spread through online social networks during crisis situations.
To approach this, we developed our first prototype simulation, modeling 100 LLM agents interacting “online.” We used GPT-4o to generate message content for each agent, which was prompted to behave as a “student.” In the simulation, agents received lockdown announcements and stay-home orders, then generated responses and posts to share within their social networks. For example, a dormitory restriction notice was interpreted by an anxious agent as a rights violation, triggering an “outrage” response that they shared on a simulated social network. Each agent was given a distinct persona generated through a two-stage prompting process, and operated with a basic state machine to guide decision-making and communication protocols. This structure enabled us to capture both individual variability and emergent patterns in how messages could be (mis)interpreted across a broad population. We implemented two subsystems, focusing on misinterpretation and propagation respectively. See the appendix (Sec A.2) for a more detailed description of the two subsystems. The simulation results showed that variations in the wording of the official announcement influenced how “student” agents exchanged messages and which agents ultimately adhered to instructions.
5.2.2. Policymaker Feedback
When we presented this system and its results to P1, P2, and P3 in October 2024, they expressed fundamental skepticism about practical utility and real-world validity. Their primary concerns centered on the difficulty of validating simulation outputs against real-world communication dynamics. For example, P3 questioned whether the model captured realistic communication patterns:
You were showing that message before with the all capital letters… there are all kinds of studies that show that people who are warned about spam emails are warned about emails having an overabundance of capital letters… I think you would find some very interesting differentiations in what the data shows.
The team then pointed out the broader challenge of validation, noting the inaccessibility of complete online communication data and the inherent unpredictability of misinformation incidents. P1 emphasized the dangers of acting on results that could not be verified:
For example, like if you implement such a simulation result, I can imagine that some high risk of the situation might be so dangerous to use such a simulation result. So maybe for especially the first stage, maybe we come up some application with low risk.
The core challenge centered around validation. Policymakers acknowledged social media misinformation as a pressing concern but emphasized the seeming impossibility of validating such simulations against ground truth data. In response, we explored alternative scenarios that both mattered to the team and offered potential points of observation and verification.
Through these discussions over the following months, we identified evacuation preparedness for the university commencement as a preferable domain. This annual event is the largest gathering on campus, and the team assumed direct responsibility for the safety of the attendees. At the same time, commencement is vulnerable to both natural hazards (e.g., severe weather) and man-made threats (e.g., targeted disruptions), making preparedness a top priority.
While such emergencies are inherently unpredictable, commencement offered a unique opportunity for observation. Because the event recurs yearly, crowd movement and communication patterns can be monitored in real time, providing data against which we might validate simulation outcomes. This made the domain both practically important and methodologically tractable, setting the stage for our second design iteration.
@@ 5.3. Second Iteration: Simulating “Commencement”
5.3.1. Simulation Goals and Design
Building on the lessons from the first iteration, we shifted focus to simulating evacuation scenarios during commencement. Unlike the previous simulation of online interactions, this iteration required modeling both spatial movement and communication among agents.
We simulated 500 LLM agents, again prompted to behave as “students,” and placed them within a rectangular space representing the university stadium where commencement is held. The simulation unfolded over multiple time steps. At the outset, agents received an official evacuation announcement. At each subsequent step, agents exchanged natural-language messages with predetermined peers, simulating online communication with their friends. They then decided which corner of the rectangle to move toward, representing exits from the stadium, with decisions generated through GPT-4o calls.
Each agent was assigned a distinct persona (based on an academic major), though they were identical in terms of physical capabilities and roles. This setup enabled us to explore how communication dynamics and individual interpretations of announcements could influence evacuation flows and crowd movement.
5.3.2. Policymaker Feedback
P1, P4, and P5 reviewed this iteration in early April 2025. While they appreciated the shift toward a scenario that could in principle be observed and validated, they immediately identified missing elements that undermined the simulation’s credibility and usefulness. For example, P4 raised practical concerns about scalability that underscored the gap between our simplified model and real-world requirements:
Is your model limited in how many agents it can handle? Commencement can be around 10,000 people, so I’m not sure how the model scales.
Across the discussion, the policymaker team pointed to several absent features: realistic physical constraints (aisles, seating sections, and bottlenecks), physical-body effects that would create realistic congestion patterns, accessibility considerations for diverse populations, the true scale of commencement attendees, the presence of volunteer coordinators that were central to their emergency procedures, and the actual announcement message used. They emphasized that without these contextual details, the simulation felt too abstract to guide concrete policy implementation.
@@ 5.4. Third Iteration: Simulating the University’s Commencement
5.4.1. Simulation Goals and Design
In response to policymakers’ feedback, we incorporated detailed stadium features into the simulation space, including concourses, aisles, and bottlenecks that reflected the actual commencement venue layout. We also increased the number of LLM agents to 3,000. Agents were implemented to adjust their movement speed dynamically based on local congestion, addressing concerns that evacuation safety and efficiency depend on crowd density. We further implemented collision-aware navigation that respected aisle and concourse boundaries while preventing unrealistic oscillations in narrow passages.
To capture variation among attendees, we explicitly introduced heterogeneous agent types. Accessibility considerations were integrated by assigning differentiated mobility parameters to agents with diverse needs. For instance, wheelchair users were placed in designated seating areas and constrained to use specific exits, mirroring actual commencement practices. In addition to “student” agents, we introduced “coordinator” agents positioned at eight fixed locations throughout the stadium. “Coordinator” agents were prompted to guide nearby “student” agents toward specific exits to improve evacuation efficiency, capturing their real-world role in emergency procedures. We consulted policymakers to obtain the official evacuation message and to identify the type of emergency they consider most important. We then used that evacuation message, as defined by the policymakers, to send to agents during a severe weather emergency—the type of emergency they believe is most likely and most concerning.
This combination of structural realism, heterogeneous mobility, and role differentiation aimed to bring the simulation closer to the practical concerns raised by policymakers.
5.4.2. Policymaker Feedback
The response to this iteration marked a significant shift in policymaker engagement. P1, P4, and P5 reviewed the system at the end of April 2025 and expressed substantially increased confidence in its potential utility. The addition of realistic scale, physical constraints and coordinator roles created what policymakers recognized as a credible representation of their operational environment.
P1 showed strong interest in the simulation output and spontaneously began envisioning how different coordinator layouts might improve evacuation efficiency, leading to concrete requests for additional scenarios:
Can you run it again until they leave the stadium? Yeah, like the whole process… This helps us figure out where to position volunteers to assist. Right now, we have four volunteers at the bottom of the stadium, but by the time the crowd reaches them, it’s too late – they’re already bumping into each other. For example, in front of the stage, maybe half could be directed right and half left. If we move the triangles (coordinators) closer to the seating sections, it would help. We may need a red triangle and three purple helper triangles for traffic, since captains alone can’t manage that volume… The real inefficiency is when students choose far exits and collide. So rather than placing captains at the exits, it may be better to position them near the sections, as that’s when people decide which way to go… This visualization is very helpful for planning locations.
P4 also immediately identified training applications that we had not originally considered:
This will be also very good for training… because we usually have slides where we explain this is your role, this is what you’re going to do, these are the locations, and if we can record a video of a scenario… just so people see this is where you can expect people.
5.4.3. Policy Implementation
This iteration marked the first concrete policy implementation. Following their review, policymakers began integrating simulation recordings into their coordinator training sessions before the 2025 commencement. They used visualization outputs to illustrate crowd dynamics and evacuation decision-making processes, replacing static slide presentations with dynamic scenario demonstrations.
Policymakers also documented our simulation approach and findings in their official after-action report following the commencement event. As the university’s authoritative record of preparedness practices, the report carries institutional weight by guiding future policies, training, and large-scale safety planning, writing: This foundational work represents an important resource in understanding crowd dynamics and optimizing emergency response strategies. The report emphasized the broader institutional value of the research: This research has been important not only for improving immediate event safety but also for informing future large-scale public safety planning. Finally, the report committed to sustained interaction of the simulation into preparedness efforts: The continued development and analysis of this simulation model will contribute to more resilient and adaptive evacuation protocols, ultimately supporting safer environments for all attendees.
@@ 5.5. Fourth Iteration: Validating Simulation with Real Commencement
5.5.1. Empirical Observation
After the fourth iteration, we had the opportunity to observe the university’s commencement alongside the policymaker team in May 2025. In preparation, we examined the physical layout of the venue before the event to understand how space was actually used during crowd movement. During the event itself, we recorded crowd movement and communication patterns with assuming anonimity under the university’s IRB confirmation. We paid particular attention to how attendees exited the stadium once the ceremony concluded. Although no emergency incidents occurred during the event, these recordings served as baseline data for our simulation models.
We also stayed in communication with the team throughout the event. For example, P1 pointed out several key locations that they were monitoring closely while attendees were leaving the stadium. These insights helped us connect our simulation assumptions with the real points of attention for practitioners, reinforcing the importance of grounding simulations in observable phenomena.
One critical finding was the central role of students’ family members. Commencement is not only about celebrating students but also about their significant others. In practice, more than half of attendees were family members. After the event ended, many students immediately sought to find and reunite with their families. As a result, the most severe congestion occurred around family seating areas, where students attempted to reach relatives, which could affect emergency evacuation. Our earlier simulations had included only “student” agents communicating with each other. This observation revealed a major gap, prompting us to reconsider simulation scale, agent roles, and validation points. Family members would need to be modeled as distinct agent types with different goals, movements, and interactions.
5.5.2. Simulation Goals and Design
Addressing the gaps identified after the May 2025 commencement observation, we developed the most comprehensive version of our system incorporating the elements policymakers had deemed missing in prior versions. This iteration scaled to nearly 13,000 agents—including both “student” and “family/friends” agents—along with 50 coordinator agents, matching the actual size and structure of the event.
Agents were designed to reflect complex social and spatial behaviors observed in the real world. “Student” and “family/friends” agents were seated in designated areas and paired (or grouped in trios), enabling them to communicate and jointly decide their next destination. Coordinators were positioned throughout the venue to guide nearby groups toward exits, mirroring real emergency procedures.
Figure 3 illustrates the overall architecture of our LLM agent simulator, including both the agent-level decision loop and the stadium-scale interface. See the Appendix (Sec A.3) for a more detailed description of the final version system  [^1] We do not consider the system, itself, to be a core contribution of this work but the design process in which the system was a technology probe..
5.5.3. Policymaker Feedback
Figure 3. Simulation overview. (A) LLM agents conceptual diagram: Agents receive inputs from three sources: Personal (e.g., accessibility needs, group membership, persona), Social (group chat, official instructions, social network), and Environmental (stadium overview, local layout, neighbor mobility). For each agent, an LLM governs decision-making and communication and hands off to a rule-based controller for embodied action. Outputs include Physical (moving/navigating, exiting), Decision (destination choice, regrouping), and Communication (ongoing group chat). (B) System display: Stadium-scale interface showing physical layouts, agents as colored dots, coordinators, and a per-agent side panel. Numbered callouts: (1) stage area; (2) students with accessibility needs; (3) seating areas for a portion of students’ family and friends; (4) seating sections for students from different majors, each major marked by a distinct color; (5) exits; (6) bleacher area with dense family-and-friends seating; (7) coordinators distributed around the field track for directing flow; (8) a per-agent side panel showing persona attributes including name, major and profile, below are group-chat messages between agents that have social relationships and within the same group. Together this system supports simulations with tens of thousands of agents (up to  13k in our largest runs).:
This iteration generated the most comprehensive validation session, held in June 2025. P1, P4, and P5 systematically compared simulation outputs against their observations from the recent commencement event and the recordings of the commencement. They confirmed that the system realistically reproduced familiar behavioral patterns and spatial dynamics, marking a substantial step forward in credibility. As shown in Fig. 4, the final simulation closely mirrored crowd dynamics observed during the actual commencement, with similar mobility patterns like congestion near the track areas close to the camera.
P4 noted how agents’ location choices mirrored authentic crowd behavior (emphasis ours):
…there are like agents who choose to stay at the boosters or… stay in the family and friends areas… on the east side and the west side… which is also what we see in the actual commencement, right? Like people around this place and this place… So this is pretty close reflection because they’re talking to each other.
P1 emphasized how this baseline validation created the foundation for trusting the system’s utility for exploring novel scenarios (emphasis ours):
Keep in mind that when we gave you all of those additional feedback pieces and you put it together when you showed a baseline evacuation without anything major, it mirrored the 20 minutes of what you all experienced with us with just the regular evacuation. And your simulation mirrored what we experience every time we have it. So that gives us some baseline trust, if you will, to see that it is connecting to what experiences we already had… so that as you all start to make additional trainings, we can take great learnings from it, but we can see how it connects and the pieces make sense as it does that. It’s not always throwing out a big red flag of oh my god something is wrong. It’s mirroring the stuff that we’ve seen before as well but then helping us identify additional pieces and so there’s a lot of value to that.
@@ 5.5.3. Policymaker Feedback
Figure 4. Comparison of crowd movement dynamics between empirical record and simulation. (A) Photograph of the 2025 commencement, showing attendee movement as the ceremony concluded. (B) Simulation snapshot of the same scenario, incorporating 13,000 agents with roles for students, family members, and coordinators, represented as colored dots. The simulation reproduced key spatial patterns observed in the real event, such as congestion on the track areas at the bottom of the frame, providing a credible basis for validating the system.:
This validation success marked another turning point: policymakers began engaging with the system not only to verify known outcomes but also to explore hypothetical emergency scenarios. They spontaneously identified policy implementation opportunities, particularly around coordinator positioning strategies and differentiated announcement messages for specific emergency types. They also expressed interest in testing how variations in coordinator layouts might affect evacuation efficiency and how tailored communication protocols could improve response effectiveness.

### v2

@@ 4.1. Iterative Design and Participants
We conducted an iterative design engagement with the policymaker team between May 2024 and August 2025, spanning approximately 16 months. This collaboration was structured around flexible meetings convened as opportunities and needs emerged, typically every 2–6 weeks during active design phases. Our goal was not to sample policymakers broadly, but to closely examine how simulation usefulness emerges through sustained engagement with a complete decision-making unit within the organization. While the broader team included many stakeholders, our design work primarily involved five core members who were directly responsible for preparedness and crisis communication: a Senior Director for Disaster Recovery and Business Continuity Services (P1), a Senior Director for Reputation and Issues Management (P2), a Director for Social Media (P3), and two Emergency Preparedness and Response Specialists (P4, P5). Together, these individuals represented both policy-level decision makers and operational specialists, allowing our process to bridge institutional strategy with on-the-ground practice.
@@
4.2. Simulation System (Final Version)
Figure 2. Simulation overview. (A) LLM agents conceptual diagram: Agents receive inputs from three sources: Personal (e.g., accessibility needs, group membership, persona), Social (group chat, official instructions, social network), and Environmental (stadium overview, local layout, neighbor mobility). For each agent, an LLM governs decision-making and communication and hands off to a rule-based controller for embodied action. Outputs include Physical (moving/navigating, exiting), Decision (destination choice, regrouping), and Communication (ongoing group chat). (B) System display: Stadium-scale interface showing physical layouts, agents as colored dots, coordinators, and a per-agent side panel. Numbered callouts: (1) stage area; (2) students with accessibility needs; (3) seating areas for a portion of students’ family and friends; (4) seating sections for students from different majors, each major marked by a distinct color; (5) exits; (6) bleacher area with dense family-and-friends seating; (7) coordinators distributed around the field track for directing flow; (8) a per-agent side panel showing persona attributes including name, major and profile, below are group-chat messages between agents that have social relationships and within the same group. Together this system supports simulations with tens of thousands of agents (up to  13k in our largest runs).
Two-panel figure summarizing our evacuation simulator. Panel A is a block diagram: three input sources—Personal (accessibility needs, group membership, persona), Social (group chat, official instructions, social network), and Environmental (stadium overview, nearby layout, neighbor mobility)—feed each agent’s LLM, which handles decision-making and communication (“Initialized” to “Discussing”, and “Discussing” to “Moving”) before handing off to a rule-based controller for embodied action (“Moving”, “Waiting,”, “Exited”). Outputs include Physical (moving and exiting), Decision (destination choice and regrouping), and Communication (group chat). Panel B is a stadium-scale display with agents rendered as dots; numbered markers indicate (1) the stage area, (2) a group of students with accessibility needs, (3) seating for some family and friends, (4) student seating sections by major (distinct colors denote different majors), (5) perimeter exits, (6) a dense bleacher area for family and friends, (7) coordinator posts around the track, and (8) a per-agent side panel showing persona fields (name, major, profile) and live group-chat messages among socially connected group members. The system supports large runs with up to approximately 13,000 agents.
Across the design process, we developed and iteratively refined an LLM-based agent simulation system. We note that the system itself is not the primary contribution of this work; nevertheless, we describe its final form to clarify the artifact through which policymakers engaged with, evaluated, and ultimately adopted simulation outputs.
Figure 2 shows the final system, which simulates approximately 13,000 agents, corresponding to the scale of the university’s commencement event. Agents represent students, family members, and emergency coordinators, each with distinct roles, goals, mobility constraints, and social relationships. Agents receive inputs from official announcements, local environmental conditions, nearby crowd dynamics, and group communication through online networks. OpenAI’s GPT-4.1 governs agents’ decision-making and message generation, while a rule-based controller handles physical movement and collision-aware navigation.
The simulation environment models the actual stadium layout, including seating sections, aisles, concourses, exits, and accessibility routes. This combination enables the system to represent both physical crowd dynamics and socially mediated behaviors, such as regrouping with family members and responding to coordinator guidance. Across scenarios, agents exhibited interpretable context-sensitive behavior: for example, they moved more urgently when informed of approaching severe weather than during routine event dispersal, and responded heterogeneously to localized bomb-threat announcements, with some agents prioritizing rapid exit while others attempted to regroup with family members or parallelized with expressing fear and uncertainty.
Policymakers primarily interacted with the system through visualizations, recorded scenario replays, and comparative scenario outputs rather than direct parameter tuning. Implementation details, including agent prompting and system architecture, are provided in the Appendix (Sec A.3).
In this paper, we treat the system as a technology probe (Hutchinson et al., 2003) — sufficiently realistic to support validation , deliberation, and preference elicitation — through how simulation usefulness emerged in practice. Thus, in the following sections, we show and analyze how usefulness emerged through engagement with this artifact, rather than evaluating the system itself.
@@ 4.3. Data Collection and Analysis
We analyzed our qualitative data as follows. The first author transcribed and coded meeting transcripts, interviews, recordings of policy implementation proposal sessions, and relevant email exchanges. Two other researchers independently reviewed the coded transcripts and provided feedback. The first author began by open coding one transcript, then discussed the emerging codes with the other coders to align on coding granularity and analytic focus. We then applied open coding across the full set of transcripts and emails, capturing both immediate design concerns and broader reflections about policy relevance (Strauss, 1987). Together, we refined the codes into higher-level categories and conducted a thematic analysis to surface recurrent patterns (Braun and Clarke, 2006). We resolved disagreements through discussion until consensus was reached. These qualitative analyses provided the foundation for the detailed accounts of our design iterations and findings presented in the following sections.
@@ 5. Design Iterations
We present the iterations in detail to show how simulation usefulness emerged through interaction with institutional constraints, rather than from technical refinement alone. Across a 16-month engagement, we developed our simulation system through multiple iterations, each shaped by policymakers’ feedback and targeted design challenges (Fig. 3). While these iterations are specific to one institution, the design challenges encountered such as validation and trust calibration reflect broader challenges in LLM agent simulations for policy use cases.
Each iteration is described only to the extent necessary to show how policymakers’ perceptions, validation criteria, and use cases evolved. Technical details are intentionally minimized.
Figure 3. Iterative design process. Across 16 months (May 2024–Aug 2025), we progressed from preparation interviews through five simulation iterations and an in-situ observation. Each phase introduced greater scale and realism (from 100 to 13,000 agents, adding roles, bottlenecks, and family dynamics) and was shaped by policymakers’ feedback. The core system used in later studies (Fig. 2) was developed by the fourth iteration. Early iterations surfaced distrust and missing realism, later iterations built credibility and training value, and by the final iteration, simulations informed adopted protocols, feasibility assessments, and official after-action reports.
A timeline showing the design process from May 2024 to August 2025. It includes preparation interviews, five iterations of simulations that grow from 100 to 13,000 agents, and an observation at a real commencement. Each stage shows how policymakers’ feedback shaped the next iteration, leading to greater realism, trust, and eventual adoption in training and protocols.
@@ 5.1. Preparation interviews: Capturing Needs and Institutional Realities
To ground the design process, we conducted three formative semi-structured interviews with emergency preparedness policymakers (P1–P3) during the first five months of the project. Interviews focused on preparedness workflows and policy development during emergencies. Insights were synthesized into stakeholder and process maps that documented key actors, responsibilities, and information flows, and were reviewed with participants. Full maps are included in the Appendix (Figs.  A2, A3, A4, and A5).
@@ 5.2. First Iteration: Simulating Communication and Misinformation Dynamics
Based on preparatory interviews, we identified online communication and misinformation as primary concerns during emergencies. P1, P2, and P3 were particularly interested in how official announcements might be misinterpreted and how such misinterpretations could spread through online networks.
To explore this space, we developed an initial prototype simulating 100 LLM agents interacting online. Agents, prompted to behave as students, received official lockdown announcements and shared generated responses through a simplified social network. Personas and lightweight state machines introduced variability in interpretation and propagation (implementation details in Appendix, Sec. A.2).
When we presented this iteration in October 2025, policymakers expressed skepticism about its practical utility. While they agreed that modeling misinformation was an important issue, their concerns centered on validation: simulated communication dynamics could not be reliably assessed against real-world behavior.
These discussions led to a key design shift: rather than pursuing simulations that focused exclusively on communication dynamics, we jointly identified evacuation preparedness for the university’s annual commencement as a more suitable domain: one that was both operationally critical and recurrent, enabling direct observation and validation.
@@ 5.3. Second Iteration: Simulating “Commencement”
Building on the limitations identified in the first iteration, we shifted focus to evacuation preparedness during the university’s annual commencement. Unlike online communication scenarios, commencement provided a recurring, institutionally significant setting in which crowd movement and response patterns could be directly observed.
We developed a simplified evacuation simulation with 500 LLM agents prompted to behave as students attending commencement. Agents were placed in an abstracted stadium layout, received an official evacuation announcement, exchanged messages with nearby peers, and selected exits over multiple time steps. This iteration deliberately prioritized observability over realism, serving as a bridge between the communication-focused prototype and a physically grounded evacuation model.
When reviewed by P1, P4, and P5 in early April 2025, policymakers immediately identified gaps that limited policy relevance. Their feedback emphasized missing contextual details essential for real evacuation planning, including realistic scale, physical constraints (aisles and bottlenecks), accessibility considerations, volunteer coordinator roles, and the use of realistic announcement language.
Rather than rejecting the simulation, policymakers treated this iteration as a diagnostic artifact. Their critiques clarified which contextual features were necessary for trust and directly motivated the next iteration’s increase in scale, environmental fidelity, and role differentiation.
@@ 5.4. Third Iteration: Simulating the University’s Commencement
Guided by policymakers’ feedback, we substantially increased the realism of the simulation to reflect the operational conditions of the university’s commencement. We incorporated key physical features they identified—including seating sections, aisles, concourses, and bottlenecks—and increased the simulation scale to 3,000 agents.
Agents reflected heterogeneous attendee roles and constraints. We introduced differentiated mobility profiles for accessibility needs and coordinator agents positioned throughout the stadium to guide evacuation, mirroring volunteer responsibilities during real emergencies. Movement speed adapted dynamically to local crowd density, enabling realistic congestion patterns to emerge.
We worked with policymakers to obtain the official evacuation message and focused on severe weather scenarios, which they identified as the most operationally relevant emergency type. Together, these changes shifted the simulation from an abstract model to one recognizable as their institutional environment.
This iteration marked a clear turning point in engagement. When P1, P4, and P5 reviewed the system in late April 2025, they expressed substantially greater confidence in its credibility and relevance. Rather than questioning validity, policymakers began using the simulation to reason about concrete planning decisions. For example, P1 used the visualization to diagnose where volunteer coordination mattered most, noting that congestion emerged well before crowds reached the exits. In their words, “by the time the crowd reaches the exits, it’s too late — they’re already bumping into each other,” whereas the key inefficiency occurred earlier, when students selected distant exits and collided along interior routes. Seeing these dynamics unfold in the simulation helped them reason about where volunteers should be positioned to intervene effectively.
Policymakers also immediately recognized the value of recorded simulations as training materials. Rather than relying on static slides, they emphasized that coordinators could “see where congestion is likely to happen and what to expect,” making the simulations a concrete way to prepare volunteers for real-world conditions. Following this review, policymakers integrated them into coordinator training sessions ahead of the 2025 commencement and documented the approach in the official after-action report. This marked the first transition of the simulation system from an exploratory technology probe (Hutchinson et al., 2003) to an institutionally-recognized tool supporting policy implementation.
@@ 5.5. Fourth Iteration: Validating Simulation with Real Commencement
Following the third iteration, we observed the university’s actual commencement alongside the policymaker team in May 2025. We examined the venue layout and recorded crowd movement as attendees exited the stadium under IRB approval. Although no emergency occurred, these observations provided baseline data for validating evacuation dynamics.
During the event, policymakers highlighted areas they routinely monitored for congestion, helping align our analytic focus with practitioner priorities. One unanticipated pattern proved especially consequential: the central role of students’ family members. More than half of attendees were family or friends, and after the ceremony many students sought to reunite with them, producing the most severe congestion near family seating areas.
Our prior simulations had modeled only student agents acting independently. This mismatch revealed a critical omission. Evacuation behavior was shaped not only by individual movement but by social regrouping, motivating a redesign that treated family members as distinct agents with different goals and interactions.
To address this gap, we developed the most comprehensive version of the system, scaling to nearly 13,000 agents—matching the actual event—and incorporating student, family/friends, and coordinator agents. Agents were seated in designated areas, grouped socially, and designed to communicate and move jointly. The stadium layout, exits, and accessibility areas reflected the real venue, enabling realistic congestion and regrouping dynamics (Fig. 2).
In June 2025, P1, P4, and P5 judged the simulation’s baseline crowd dynamics to align with their routine commencement experience (Fig. 4), which established sufficient trust to explore hypothetical emergencies.
@@
Figure 4. Comparison of crowd movement dynamics between empirical observation and simulation. (A) Photograph of the 2025 commencement, showing attendee movement as the ceremony concluded under routine (non-emergency) conditions. (B) Snapshot from the corresponding simulation, incorporating 13,000 agents with roles for students, family members, and coordinators, represented (colored dots). The simulation (Fig. 2) reproduced key spatial patterns observed in the event, including congestion in the track areas at the bottom of the frame, supporting alignment between observed and simulated crowd dynamics.
Two side-by-side panels compare crowd movement. The left panel is a photograph of the university’s 2025 commencement, showing large groups of people exiting seating sections and gathering on the track. The right panel is a simulation diagram of the same venue with thousands of small dots representing agents. Both images highlight similar congestion patterns near the track areas close to the camera, illustrating how simulated dynamics align with real-world observations.